OpenStudy (anonymous):

find dr/du, dr/dv, dr/dt r=xlny x=3u+vt y=uvt

OpenStudy (anonymous):

I dont understand the question

OpenStudy (anonymous):

It may just be me though.

OpenStudy (anonymous):

find \[\delta r/\delta u\] \[\delta r/\delta v\] \[\delta r/\delta t\]

OpenStudy (anonymous):

does that make it more understandable?

OpenStudy (anonymous):

Nope! probably cause its out of my league :D Haven't learned that yet.

OpenStudy (anonymous):

haha thanks for trying, how about you @KingGeorge ?

OpenStudy (anonymous):

What subject is that? Just curious:) I'm gonna end up taking it im sure

OpenStudy (anonymous):

its like.. calc 3? i think.. it depends on where you're taking it.. for my its either math 252, or calc 20c, but i think the general subject is calc 3, specifically, multivariable differential equations, and partial fraction chain rule decomposition

OpenStudy (anonymous):

Yup I need calc 3 for my comp science major. Im only in trig right now.

OpenStudy (anonymous):

good luck to you! i dunno about you but i'm terrriibbblleee at math, and i have a really bad prof, so i'm struggling to get a D -_-"

OpenStudy (anonymous):

Its been my strong subject I guess. I'm almost done with my trig class and I have 90,97,98 test grades so far. But yea, Professors make the biggest difference.

OpenStudy (anonymous):

I have the ones who teach straight from a powerpoint word by word..

OpenStudy (anonymous):

for most topics it's not that bad, for math its terrible.. she lectures straight out of the book.. doesn't teach anything conceptual, just tells you how to do stuff, and then doesn't really let us ask questions

OpenStudy (anonymous):

Thats dumb haha

OpenStudy (dumbcow):

first get r in terms of u,v,t then take derivative with respect to each variable, holding the rest as constants r=(3u+vt) ln(uvt) i'll do dr/du, v and t will act as constants use product rule dr/du = 3ln(uvt) + (3u+vt)(1/u)

OpenStudy (anonymous):

You're a smart cow!:D

OpenStudy (anonymous):

let me tell you what i got, cuz it's wrong, and maybe you could see what i'm doing wrong? for dr/du dr/dx = lny dy/du = 3 dr/dy = x/y dx/du = xvt 3ln(uvt) + (3u+vt)/v

OpenStudy (anonymous):

and, this is in the chapter of the chain rule, so i think i have to use it haha

OpenStudy (anonymous):

@eliassaab i'm not quite sure how you got o.O

OpenStudy (dumbcow):

oh i see, well you get same answer how do you end up with a v on bottom?

OpenStudy (dumbcow):

dx/du = 3 dy/du = vt

OpenStudy (anonymous):

so using the formula dr/du= (dr/dx)(dx/du) + (dr/dy)(dy/du) dr/dy = x/y so plugging that in you get [(3u+vt) / uvt] x [vt] so the vt on the bottom cancels, i think?

OpenStudy (anonymous):

sorry dy/du is vt

OpenStudy (dumbcow):

correct, leaving u on the bottom

OpenStudy (anonymous):

sorry that was a typo!! -_- i have a hard time telling my u's and v's apart... 3ln(uvt) + (3u+vt)/u

OpenStudy (dumbcow):

so did we get the same wrong answer ?

OpenStudy (anonymous):

the right answer is 3ln(uvt)+3+vt/u and i'm not sure whats different, what i did wrong

OpenStudy (dumbcow):

no we are correct, they just split the fraction up (3u+vt)/u = 3 + vt/u

OpenStudy (anonymous):

but there's a u at the top in the first part, and there sin't one in the second.. i'm so confused

OpenStudy (dumbcow):

eliasaab, how are you obtaining this equation by substitution of x,y r=lmt2uv2+3lmtu2v

OpenStudy (dumbcow):

itz, ?? 3u/u = 3

OpenStudy (anonymous):

but.. its part of the equation.. are you splitting it like 3u/u + vt/u ?! you can do that?

OpenStudy (anonymous):

Replacing x and y in r, we obtain \[ r=3 u \ln (t u v)+t v \ln (t u v) \] \[ \frac {\partial r }{du}=\frac{t v}{u}+3 \ln (t u v)+3 \] \[ \frac {\partial r }{dv}=t \ln (t u v)+t+\frac{3 u}{v} \] \[ \frac {\partial r }{dt}=v \ln (t u v)+\frac{3 u}{t}+v \]

OpenStudy (anonymous):

could you explain how you did the derivations for that? if you don't mind?

OpenStudy (anonymous):

You have r as a function of u , v and t, You take the partial derivatives of r with respect to u, then to v, then to t